OpenStudy (anonymous):

find the derivative: (x-1)e^3x+2

OpenStudy (lalaly):

you use the product rule let f = (x-1)e^3x+2 u=x-1 v=e^3x \[\large{f'=u'v + v'u}+\frac{d}{dx}2\]\[\large{f'= (1 \times e^{3x}) +(x-1)\times 3e^{3x} +0}\]

OpenStudy (anonymous):

this is what I got......